OpenStudy (anonymous):

Alton has been recording the number of apples on his apple tree each week: Week Apples 2 15 3 22 4 29 5 36 The function representing Alton's apples per week is f(w) = 7w + 1. What does the 1 represent? The number of apples at the start The week number when Alton recorded the apples The number of apples Alton had in total The number the apples increased by each week

OpenStudy (solomonzelman):

y=mx+b, where b is the y-intercept What does it mean when w=0 ?

OpenStudy (solomonzelman):

(w is the number of weeks, and f(w) is the number of apples after w weeks )

OpenStudy (anonymous):

ok so when w =0 that means its 8? sorry just a little confused

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(w) = 7w + 1 }\) plugging in 0 for w, \(\normalsize\color{blue}{ f(0) = 7(0) + 1 }\) \(\normalsize\color{blue}{ f(0) = 0+ 1 }\) \(\normalsize\color{blue}{ f(0) = ? }\) you tell me

OpenStudy (anonymous):

oh ok so it=1

OpenStudy (solomonzelman):

I love disconnecting from the site ! ... -:( Anyway, \(\normalsize\color{blue}{ f(w) = 7w+\color{red}{1} }\) (this 1 ) represents the number of apples when w=0, or the number of apples to begin with.

OpenStudy (anonymous):

yeah i hate hen that happens. ok so then that means the answer would be The number of apples at the start

OpenStudy (solomonzelman):

Yes, exactly

OpenStudy (anonymous):

yayy ok thanks for your help! (:

OpenStudy (solomonzelman):

Anytime!